Yahaira:

x−2.6y=−4.8 −x+3.4y=7.2

Yahaira:

Can someone help me solve this pls?

CuteTeddyUwU:

is it two separate equations or one- ?

Yahaira:

Well my teacher said you can't solve one without the other

ZombieLover:

What are we trying to solve?

Yahaira:

x−2.6y=−4.8 −x+3.4y=7.2 The x and y

ZombieLover:

so top equation first then we adding on the second one together or? adding both sum